OpenStudy (anonymous):

What is the slope of the line that contains the points (-2, 9) and (4, 7)?

jimthompson5910 (jim_thompson5910):

Use the slope formula m = (y2-y1)/(x2-x1) m = ( 7- 9)/(4-(-2)) m = ( 7- 9)/(4+2) I'll let you finish

OpenStudy (anonymous):

Thank you!